# GenIAS: Generator for Instantiating Anomalies in time Series

### v1

@@ G.2.1. Critical Difference diagrams
Multivariate Univariate Metric Cat. Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg Aff Classic OCSVM (1999) 0.742 0.761 0.733 0.197 0.425 0.711 0.595 0.923 0.761 0.710 0.798 0.663 LOF (2000) 0.775 0.762 0.666 0.613 0.227 0.745 0.631 0.906 0.674 0.718 0.766 0.676 IForest (2008) 0.719 0.728 0.699 0.535 0.259 0.824 0.627 0.907 0.760 0.686 0.784 0.680 PU LSTM-VAE (2018) 0.824 0.782 0.630 0.324 0.312 0.726 0.600 M M M M M US-WI Donut (2018) U U U U U U U 0.777 0.682 0.685 0.715 - Omni (2019) 0.680 0.779 0.706 0.156 0.230 0.608 0.527 M M M M M THOC (2020) 0.778 0.781 0.677 0.679 0.321 0.713 0.658 0.796 0.676 0.642 0.705 0.674 MTAD-GAT(2020) 0.833 0.848 0.731 0.355 0.308 0.745 0.637 M M M M M AnomalyTran (2021) 0.774 0.766 0.664 0.482 0.551 0.537 0.629 M M M M M TranAD (2022) 0.771 0.751 0.719 0.606 0.233 0.661 0.624 0.917 0.779 0.698 0.798 0.682 TS2Vec (2022) 0.694 0.695 0.750 0.700 0.520 0.728 0.681 0.867 0.818 0.671 0.785 0.716 TimesNet (2023) 0.769 0.710 0.795 0.693 0.829 0.710 0.751 0.917 0.844 0.712 0.824 0.775 Dcdetector (2023) 0.706 0.680 0.528 0.693 0.671 0.710 0.665 0.737 0.666 0.671 0.691 0.674 US-I NCAD (2022) 0.685 0.664 0.732 0.693 0.887 0.710 0.729 0.645 0.753 0.747 0.715 0.724 CutAddPaste (2024) 0.762 0.753 0.693 0.764 0.659 0.875 0.751 0.832 0.818 0.820 0.823 0.775 COUTA (2024) 0.785 0.762 0.764 0.645 0.592 0.731 0.713 0.906 0.851 0.730 0.829 0.752 CARLA (2025) 0.816 0.808 0.733 0.592 0.756 0.932 0.773 0.874 0.701 0.805 0.793 0.780 Ours GenIAS 0.851 0.856 0.731 0.596 0.763 0.935 0.789 0.921 0.788 0.814 0.841 0.806 PATE Classic OCSVM (1999) 0.267 0.270 0.363 0.698 0.191 0.704 0.416 0.780 0.343 0.033 0.385 0.405 LOF (2000) 0.242 0.251 0.319 0.187 0.042 0.706 0.291 0.690 0.159 0.056 0.302 0.295 IForest (2008) 0.195 0.197 0.306 0.743 0.191 0.799 0.405 0.735 0.325 0.027 0.362 0.391 PU LSTM-VAE (2018) 0.318 0.278 0.436 0.693 0.061 0.753 0.423 M M M M M US-WI Donut (2018) U U U U U U U 0.358 0.120 0.018 0.165 - Omni (2019) 0.143 0.162 0.422 0.702 0.076 0.709 0.369 M M M M M THOC (2020) 0.307 0.218 0.147 0.508 0.104 0.563 0.308 0.488 0.317 0.016 0.274 0.296 MTAD-GAT(2020) 0.391 0.371 0.445 0.115 0.247 0.707 0.379 M M M M M AnomalyTran (2021) 0.267 0.283 0.332 0.683 0.112 0.685 0.394 M M M M M TranAD (2022) 0.238 0.279 0.458 0.192 0.141 0.688 0.333 0.777 0.408 0.025 0.403 0.356 TS2Vec (2022) 0.143 0.150 0.137 0.147 0.055 0.530 0.194 0.636 0.247 0.144 0.342 0.243 TimesNet (2023) 0.282 0.207 0.452 0.083 0.240 0.636 0.317 0.760 0.355 0.042 0.386 0.340 Dcdetector (2023) 0.176 0.135 0.097 0.124 0.025 0.455 0.169 0.184 0.052 0.025 0.087 0.141 US-I NCAD (2022) 0.157 0.161 0.129 0.107 0.272 0.556 0.230 0.145 0.173 0.130 0.149 0.203 CutAddPaste (2024) 0.269 0.347 0.200 0.613 0.252 0.842 0.421 0.237 0.741 0.262 0.413 0.418 COUTA (2024) 0.270 0.244 0.447 0.159 0.285 0.756 0.360 0.684 0.335 0.054 0.358 0.359 CARLA (2025) 0.493 0.413 0.488 0.655 0.193 0.911 0.526 0.695 0.378 0.301 0.458 0.503 Ours GenIAS 0.571 0.516 0.553 0.718 0.274 0.941 0.596 0.838 0.456 0.331 0.542 0.578 AUC-ROC Classic OCSVM (1999) 0.605 0.634 0.762 0.799 0.446 0.714 0.660 0.918 0.728 0.567 0.738 0.686 LOF (2000) 0.610 0.584 0.708 0.660 0.370 0.789 0.620 0.890 0.604 0.575 0.690 0.643 IForest (2008) 0.571 0.579 0.768 0.842 0.927 0.862 0.758 0.915 0.749 0.570 0.745 0.754 PU LSTM-VAE (2018) 0.618 0.650 0.771 0.806 0.506 0.763 0.686 M M M M M US-WI Donut (2018) U U U U U U U 0.676 0.640 0.498 0.605 - Omni (2019) 0.387 0.570 0.794 0.811 0.430 0.696 0.615 M M M M M THOC (2020) 0.654 0.604 0.678 0.318 0.395 0.578 0.538 0.873 0.799 0.558 0.743 0.606 MTAD-GAT(2020) 0.674 0.641 0.812 0.822 0.590 0.749 0.715 M M M M M AnomalyTran (2021) 0.641 0.653 0.764 0.809 0.698 0.678 0.707 M M M M M TranAD (2022) 0.632 0.645 0.808 0.677 0.519 0.677 0.660 0.889 0.763 0.515 0.722 0.681 TS2Vec (2022) 0.503 0.517 0.496 0.548 0.501 0.573 0.523 0.698 0.567 0.501 0.589 0.545 TimesNet (2023) 0.664 0.578 0.837 0.243 0.924 0.598 0.641 0.892 0.750 0.597 0.746 0.676 Dcdetector (2023) 0.535 0.519 0.579 0.504 0.504 0.499 0.523 0.603 0.504 0.509 0.539 0.528 US-I NCAD (2022) 0.491 0.493 0.572 0.365 0.692 0.592 0.534 0.567 0.671 0.609 0.616 0.561 CutAddPaste (2024) 0.673 0.631 0.631 0.777 0.789 0.464 0.661 0.684 0.810 0.617 0.704 0.675 COUTA (2024) 0.661 0.606 0.811 0.643 0.660 0.802 0.697 0.911 0.787 0.559 0.752 0.716 CARLA (2025) 0.694 0.677 0.657 0.783 0.696 0.510 0.670 0.543 0.604 0.735 0.627 0.655 Ours GenIAS 0.745 0.718 0.633 0.786 0.811 0.536 0.705 0.778 0.691 0.778 0.749 0.720
(a) KPI - Entity c69a50cf-ee03-3bd7-831e-407d36c7ee91
(b) SMD - Entity machine-2-3
(c) UCR - Entity 222_UCR_Anomaly_mit14046longtermecg_56123_91200_91700
@@
Figure 10. Each plot visualises a different time series (TS) injection or generation method using t-SNE. Blue markers represent the original anomalous TS, while orange markers represent the injected or generated anomalous TS.
G.2.3. Additional Visualizations

### v2

@@ G.2.1 Critical Difference diagrams
Multivariate Univariate Metric Cat. Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg AUPR Classic OCSVM (1999) 0.202 0.256 0.327 0.696 0.163 0.608 0.375 0.703 0.300 0.028 0.344 0.365 LOF (2000) 0.217 0.248 0.277 0.185 0.031 0.624 0.264 0.615 0.109 0.043 0.256 0.261 IForest (2008) 0.182 0.190 0.278 0.739 0.162 0.739 0.382 0.614 0.288 0.021 0.308 0.357 PU LSTM-VAE (2018) 0.285 0.258 0.395 0.685 0.045 0.674 0.390 M M M M M US-WI Donut (2018) U U U U U U U 0.264 0.078 0.015 0.119 U Omni (2019) 0.149 0.115 0.365 0.713 0.034 0.587 0.327 M M M M M THOC (2020) 0.240 0.195 0.107 0.537 0.076 0.432 0.265 0.349 0.229 0.014 0.197 0.242 MTAD-GAT(2020) 0.335 0.339 0.401 0.095 0.200 0.613 0.331 M M M M M AnomalyTran (2021) 0.236 0.264 0.273 0.680 0.059 0.568 0.347 M M M M M TranAD (2022) 0.273 0.287 0.412 0.192 0.119 0.586 0.312 0.691 0.285 0.022 0.333 0.319 TS2Vec (2022) 0.132 0.148 0.113 0.136 0.040 0.408 0.163 0.491 0.221 0.009 0.240 0.189 TimesNet (2023) 0.283 0.208 0.385 0.082 0.189 0.508 0.276 0.671 0.237 0.032 0.313 0.288 Dcdetector (2023) 0.129 0.124 0.043 0.126 0.011 0.326 0.127 0.041 0.018 0.009 0.023 0.092 US-I NCAD (2022) 0.146 0.156 0.096 0.106 0.135 0.407 0.174 0.067 0.089 0.103 0.086 0.145 CutAddPaste (2024) 0.242 0.322 0.153 0.582 0.159 0.808 0.378 0.116 0.602 0.188 0.302 0.352 COUTA (2024) 0.247 0.248 0.400 0.159 0.264 0.685 0.334 0.555 0.282 0.049 0.295 0.321 CARLA (2025) 0.501 0.448 0.507 0.681 0.201 0.814 0.525 0.627 0.299 0.247 0.391 0.481 Ours GenIAS+CARLA 0.529 0.467 0.512 0.704 0.236 0.882 0.555 0.794 0.357 0.280 0.477 0.529 Aff Classic OCSVM (1999) 0.742 0.761 0.733 0.197 0.425 0.711 0.595 0.923 0.761 0.710 0.798 0.663 LOF (2000) 0.775 0.762 0.666 0.613 0.227 0.745 0.631 0.906 0.674 0.718 0.766 0.676 IForest (2008) 0.719 0.728 0.699 0.535 0.259 0.824 0.627 0.907 0.760 0.686 0.784 0.680 PU LSTM-VAE (2018) 0.824 0.782 0.630 0.324 0.312 0.726 0.600 M M M M M US-WI Donut (2018) U U U U U U U 0.777 0.682 0.685 0.715 - Omni (2019) 0.680 0.779 0.706 0.156 0.230 0.608 0.527 M M M M M THOC (2020) 0.778 0.781 0.677 0.679 0.321 0.713 0.658 0.796 0.676 0.642 0.705 0.674 MTAD-GAT(2020) 0.833 0.848 0.731 0.355 0.308 0.745 0.637 M M M M M AnomalyTran (2021) 0.774 0.766 0.664 0.482 0.551 0.537 0.629 M M M M M TranAD (2022) 0.771 0.751 0.719 0.606 0.233 0.661 0.624 0.917 0.779 0.698 0.798 0.682 TS2Vec (2022) 0.694 0.695 0.750 0.700 0.520 0.728 0.681 0.867 0.818 0.671 0.785 0.716 TimesNet (2023) 0.769 0.710 0.795 0.693 0.829 0.710 0.751 0.917 0.844 0.712 0.824 0.775 Dcdetector (2023) 0.706 0.680 0.528 0.693 0.671 0.710 0.665 0.737 0.666 0.671 0.691 0.674 US-I NCAD (2022) 0.685 0.664 0.732 0.693 0.887 0.710 0.729 0.645 0.753 0.747 0.715 0.724 CutAddPaste (2024) 0.762 0.753 0.693 0.764 0.659 0.875 0.751 0.832 0.818 0.820 0.823 0.775 COUTA (2024) 0.785 0.762 0.764 0.645 0.592 0.731 0.713 0.906 0.851 0.730 0.829 0.752 CARLA (2025) 0.816 0.808 0.733 0.592 0.756 0.932 0.773 0.874 0.701 0.805 0.793 0.780 Ours GenIAS+CARLA 0.851 0.856 0.731 0.596 0.763 0.935 0.789 0.921 0.788 0.814 0.841 0.806 PATE Classic OCSVM (1999) 0.267 0.270 0.363 0.698 0.191 0.704 0.416 0.780 0.343 0.033 0.385 0.405 LOF (2000) 0.242 0.251 0.319 0.187 0.042 0.706 0.291 0.690 0.159 0.056 0.302 0.295 IForest (2008) 0.195 0.197 0.306 0.743 0.191 0.799 0.405 0.735 0.325 0.027 0.362 0.391 PU LSTM-VAE (2018) 0.318 0.278 0.436 0.693 0.061 0.753 0.423 M M M M M US-WI Donut (2018) U U U U U U U 0.358 0.120 0.018 0.165 - Omni (2019) 0.143 0.162 0.422 0.702 0.076 0.709 0.369 M M M M M THOC (2020) 0.307 0.218 0.147 0.508 0.104 0.563 0.308 0.488 0.317 0.016 0.274 0.296 MTAD-GAT(2020) 0.391 0.371 0.445 0.115 0.247 0.707 0.379 M M M M M AnomalyTran (2021) 0.267 0.283 0.332 0.683 0.112 0.685 0.394 M M M M M TranAD (2022) 0.238 0.279 0.458 0.192 0.141 0.688 0.333 0.777 0.408 0.025 0.403 0.356 TS2Vec (2022) 0.143 0.150 0.137 0.147 0.055 0.530 0.194 0.636 0.247 0.144 0.342 0.243 TimesNet (2023) 0.282 0.207 0.452 0.083 0.240 0.636 0.317 0.760 0.355 0.042 0.386 0.340 Dcdetector (2023) 0.176 0.135 0.097 0.124 0.025 0.455 0.169 0.184 0.052 0.025 0.087 0.141 US-I NCAD (2022) 0.157 0.161 0.129 0.107 0.272 0.556 0.230 0.145 0.173 0.130 0.149 0.203 CutAddPaste (2024) 0.269 0.347 0.200 0.613 0.252 0.842 0.421 0.237 0.741 0.262 0.413 0.418 COUTA (2024) 0.270 0.244 0.447 0.159 0.285 0.756 0.360 0.684 0.335 0.054 0.358 0.359 CARLA (2025) 0.493 0.413 0.488 0.655 0.193 0.911 0.526 0.695 0.378 0.301 0.458 0.503 Ours GenIAS+CARLA 0.571 0.516 0.553 0.718 0.274 0.941 0.596 0.838 0.456 0.331 0.542 0.578 AUC-ROC Classic OCSVM (1999) 0.605 0.634 0.762 0.799 0.446 0.714 0.660 0.918 0.728 0.567 0.738 0.686 LOF (2000) 0.610 0.584 0.708 0.660 0.370 0.789 0.620 0.890 0.604 0.575 0.690 0.643 IForest (2008) 0.571 0.579 0.768 0.842 0.927 0.862 0.758 0.915 0.749 0.570 0.745 0.754 PU LSTM-VAE (2018) 0.618 0.650 0.771 0.806 0.506 0.763 0.686 M M M M M US-WI Donut (2018) U U U U U U U 0.676 0.640 0.498 0.605 - Omni (2019) 0.387 0.570 0.794 0.811 0.430 0.696 0.615 M M M M M THOC (2020) 0.654 0.604 0.678 0.318 0.395 0.578 0.538 0.873 0.799 0.558 0.743 0.606 MTAD-GAT(2020) 0.674 0.641 0.812 0.822 0.590 0.749 0.715 M M M M M AnomalyTran (2021) 0.641 0.653 0.764 0.809 0.698 0.678 0.707 M M M M M TranAD (2022) 0.632 0.645 0.808 0.677 0.519 0.677 0.660 0.889 0.763 0.515 0.722 0.681 TS2Vec (2022) 0.503 0.517 0.496 0.548 0.501 0.573 0.523 0.698 0.567 0.501 0.589 0.545 TimesNet (2023) 0.664 0.578 0.837 0.243 0.924 0.598 0.641 0.892 0.750 0.597 0.746 0.676 Dcdetector (2023) 0.535 0.519 0.579 0.504 0.504 0.499 0.523 0.603 0.504 0.509 0.539 0.528 US-I NCAD (2022) 0.491 0.493 0.572 0.365 0.692 0.592 0.534 0.567 0.671 0.609 0.616 0.561 CutAddPaste (2024) 0.673 0.631 0.631 0.777 0.789 0.464 0.661 0.684 0.810 0.617 0.704 0.675 COUTA (2024) 0.661 0.606 0.811 0.643 0.660 0.802 0.697 0.911 0.787 0.559 0.752 0.716 CARLA (2025) 0.694 0.677 0.657 0.783 0.696 0.510 0.670 0.543 0.604 0.735 0.627 0.655 Ours GenIAS+CARLA 0.745 0.718 0.633 0.786 0.811 0.536 0.705 0.778 0.691 0.778 0.749 0.720
(a) SMAP - Entity E5
(b) KPI - Entity c69a50cf-ee03-3bd7-831e-407d36c7ee91
(c) SMD - Entity machine-2-3
(d) UCR - Entity 222_UCR_Anomaly_mit14046longtermecg_56123_91200_91700
@@
Figure 10: Each plot visualizes a different time series (TS) injection or generation method using t-SNE. Blue markers represent the original anomalous TS, while orange markers represent the injected or generated anomalous TS.
G.2.3 Additional Visualizations